Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for oneCan you correct the wrong inputs please
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (2.50-2.40)*13,800

= 1380 F

.

Materials Quantity Variance = (SQ-AQ) *SP

= (1380*7.10 - 12,420)*2.50

= 6555 U